# TrimCaching: Parameter-sharing Edge Caching for AI Model Downloading

### v1

@@ V-A Successive Greedy Algorithm and the Constant-approximation Guarantee
Suppose that each sub-problem can be solved optimally. In the considered special case, the TrimCaching Spec algorithm can obtain a solution $\hat{{\bf{X}}}$ for $\mathcal{P}1.1$ which is lower bounded by $U\left(\hat{{\bf{X}}}\right)\geq\frac{1}{2}U\left({\bf{X}}^{*}\right)$, where ${\bf{X}}^{*}$ is the optimal solution of $\mathcal{P}1.1$.
@@
One condition to achieve the constant-approximation guarantee of the TrimCaching Spec algorithm is obtaining the optimal solution to $\mathcal{P}2.1_{m}$. In the next subsection, we will introduce how to solve it in polynomial time.
@@ V-B DP-based Approach and the $\epsilon$-optimal Solution
We propose a DP-based algorithm to solve $\mathcal{P}2.1_{m}$. The submodularity of the constraint in $\mathcal{P}2.1_{m}$ results from the shared parameter blocks among models. To address the submodular constraints, our core idea is to make the caching decisions for shared parameter blocks and specific parameter blocks separately. We use $\mathcal{A}$ to denote a set where each element represents a combination of shared parameter blocks. All the possible combinations are included in $\mathcal{A}$. For example, $\{1,2,3,4,5,15\}$ and $\{10,12,13,15\}$ are two elements in $\mathcal{A}$ for Fig. 4. For a combination $\mathcal{N}$, all models include $\mathcal{N}$ is denoted by $\mathcal{I}_{\mathcal{N}}$. For example, if $\mathcal{N}$ is $\{1,2,3,4,5,15\}$, $\mathcal{I}_{\mathcal{N}}$ is the set of model 1 and 2 for Fig. 4. Additionally, we use $d_{\mathcal{N}}$ to denote the occupied capacity of edge server $m$ for the shared parameter block combination $\mathcal{N}$. In the above example, $d_{\mathcal{N}}$ is the total size of parameter blocks 1, 2, 3, 4, 5, 15.
@@
The process of updating $\mathcal{T}\left(e_{\mathcal{N}},w_{\mathcal{N}}\right)$ for edge server $m$ is shown in Fig. 5. Specifically, if $w_{\mathcal{N}}\geq\frac{u\left(m,e_{\mathcal{N}}\right)}{\delta_{m,\mathcal{N}}}$, i.e., the number of cache hits of model $e_{\mathcal{N}}$ is smaller than the $w_{\mathcal{N}}$-th value of the number of cache hits, $\mathcal{T}\left(e_{\mathcal{N}},w_{\mathcal{N}}\right)$ is the minimum value between $\mathcal{T}\left(e_{\mathcal{N}}-1,w_{\mathcal{N}}\right)$ and $\mathcal{T}\left(e_{\mathcal{N}}-1,w_{\mathcal{N}}-\frac{u\left(m,e_{\mathcal{N}}\right)}{\delta_{m,\mathcal{N}}}\right)+D_{\mathcal{N}}\left(e_{\mathcal{N}}\right)$. If $\mathcal{T}\left(e_{\mathcal{N}}-1,w_{\mathcal{N}}\right)$ is smaller, it means that a combination of certain models in the first $e_{\mathcal{N}}-1$ models can achieve the $w_{\mathcal{N}}$-th value of the number of cache hits with a smaller data size. Thus, it is also the data size for the first $e_{\mathcal{N}}$ models to achieve the $w_{\mathcal{N}}$-th value of the number of cache hits. If $\mathcal{T}\left(e_{\mathcal{N}}-1,w_{\mathcal{N}}-\frac{u\left(m,e_{\mathcal{N}}\right)}{\delta_{m,\mathcal{N}}}\right)+D_{\mathcal{N}}\left(e_{\mathcal{N}}\right)$ is smaller, it represents the $w_{\mathcal{N}}$-th value of the number of cache hits $w_{\mathcal{N}}\delta_{m,\mathcal{N}}$ is the sum of the number of cache hits $w_{\mathcal{N}}\delta_{m,\mathcal{N}}-u\left(m,e_{\mathcal{N}}\right)$ of some models in the first $e_{\mathcal{N}}-1$ models and the number of cache hits $u\left(m,e_{\mathcal{N}}\right)$ of model $e_{\mathcal{N}}$, thus adding $D_{\mathcal{N}}\left(e_{\mathcal{N}}\right)$ into $\mathcal{T}\left(e_{\mathcal{N}}-1,w_{\mathcal{N}}-\frac{u\left(m,e_{\mathcal{N}}\right)}{\delta_{m,\mathcal{N}}}\right)$. On the other hand, if $w_{\mathcal{N}}<\frac{u\left(m,e_{\mathcal{N}}\right)}{\delta_{m,\mathcal{N}}}$, i.e., the number of cache hits of model $e_{\mathcal{N}}$ is already greater than the $w_{\mathcal{N}}$-th value of the number of cache hits, the $w_{\mathcal{N}}$-th value of the number of cache hits can only be realized by certain models in the first $e_{\mathcal{N}}-1$ models. Finally, after updating all $\mathcal{T}\left(e_{\mathcal{N}},w_{\mathcal{N}}\right)$, the maximum expected number of cache hits of edge server $m$ with $\mathcal{N}$, i.e., $w^{*}_{\mathcal{N}}\delta_{m,\mathcal{N}}$, can be determined, where
@@
where ^$\hat{\mathcal{I}}_{m}=\left\{i\mid\hat{x}_{m,i}\in\hat{{\bf{X}}}_{m}\right\}$.
@@
To further accelerate the computation of (18), we propose a rounding method to balance precision and computational efficiency. When determining the maximum expected number of cache hits of $\mathcal{I}_{\mathcal{N}}$, a smaller value of $\delta_{m,\mathcal{N}}$ leads to a larger value of $\frac{\sum\limits_{i\in\mathcal{I}_{\mathcal{N}}}u\left(m,i\right)}{\delta_{m,\mathcal{N}}}+1$ and more number of feasible $w_{\mathcal{N}}$, further increasing the complexity. To facilitate the execution, we define a factor $\epsilon$. Then, we round $u\left(m,i\right)$ into
@@
where $u_{m,\min}=\mathop{\min}\limits_{i\in\mathcal{I}}u\left(m,i\right)$. $\epsilon$ is a constant and $\epsilon\in\left(0,1\right]$. As a result, when determining the maximum expected cache hit ratio for edge server $m$, some $u\left(m,i\right)$ of models in $\mathcal{I}_{\mathcal{N}}$ can be rounded into the same value, and the number of feasible values of $w_{\mathcal{N}}$ can be decreased, thereby accelerating the DP process for $\mathcal{I}_{\mathcal{N}}$. In the rounding process, we denote $\dot{\delta}_{m,\mathcal{N}}$ by the granularity of $\dot{u}\left(m,i\right)$ for models in $\mathcal{I}_{\mathcal{N}}$. Similarly, $\mathcal{T}\left(e_{\mathcal{N}},\dot{w}_{\mathcal{N}}\right)$ is the smallest data size of specific parameter blocks that have to be cached on edge server $m$ for achieving the $\dot{w}_{\mathcal{N}}$-th value of the number of rounded cache hits with the first $e_{\mathcal{N}}$ models in $\mathcal{I}_{\mathcal{N}}$. Besides, we define $\dot{w}^{*}_{\mathcal{N}}=\mathop{\arg\max}\limits_{\dot{w}_{\mathcal{N}}}\{\dot{w}_{\mathcal{N}}\mid\mathcal{T}\left(\left|\mathcal{I}_{\mathcal{N}}\right|,\dot{w}_{\mathcal{N}}\right)\leq Q_{m}-d_{\mathcal{N}}\}$ and $\dot{w}^{*}_{\mathcal{N}^{*}}\delta_{m,\mathcal{N}^{*}}=\mathop{\max}\limits_{\mathcal{N}}\{\dot{w}^{*}_{\mathcal{N}}\delta_{m,\mathcal{N}}\}$. At last, in $\mathcal{P}2.1_{m}$, the model caching decision corresponding to $\dot{w}^{*}_{\mathcal{N}^{*}}$ is denoted by $\dot{{\bf{X}}}_{m}$, and
@@
The complexity of the TrimCaching Spec algorithm is $O\left(MI\right)$, which means that the TrimCaching Spec algorithm can execute in polynomial time.
@@
The TrimCaching Spec algorithm can obtain a solution $\hat{{\bf{X}}}$ for $\mathcal{P}1.1$ which is lower bounded by $U\left(\hat{{\bf{X}}}\right)\geq\frac{1-\epsilon}{2}U\left({\bf{X}}^{*}\right)$, if the TrimCaching Spec algorithm adopts Algorithm 3 to obtain an $\epsilon$-optimal solution for each sub-problem.
@@ VI The General Case: Arbitrary Parameter Sharing
This section examines the general case of $\mathcal{P}1.1$, where models can arbitrarily share parameter blocks, unlike the special case with a small fixed number of shared parameter blocks. Formally speaking, the number of shared parameter blocks among models may increase with the scale of the model library, such that searching for all combinations of shared parameter blocks results in an exponential time complexity in terms of the problem scale, which should be avoided.
Our proposed greedy algorithm for $\mathcal{P}1.1$ is elaborated as follows and outlined in Algorithm 4. The key part of the TrimCaching Gen algorithm is in line 4. In the $l$-th step, the algorithm first calculates the increased cache hit ratio for each $x_{m,i}$ based on the previous placement decision ${\bf{X}}^{l-1}$ made in the $l-1$-th step. Subsequently, the TrimCaching Gen algorithm identifies $\{m^{*},i^{*}\}$ which yields the maximum increased cache hit ratio, i.e., $U\left({\bf{X}}^{l-1}\cup\left\{x_{m^{*},i^{*}}\right\}\right)-U\left({\bf{X}}^{l-1}\right)$, while ensuring that $g_{m}\left({\bf{X}}_{m}^{l-1}\cup\{x_{m^{*},i^{*}}\}\right)$ does not exceed the edge server capacity $Q_{m}$, where ${\bf{X}}_{m}^{l-1}$ is the model placement result of edge server $m$ in the $l-1$-th step. $x_{m^{*},i^{*}}$ will be added into ${\bf{X}}^{l-1}$, yielding ${\bf{X}}^{l}={\bf{X}}^{l-1}\cup\left\{x_{m^{*},i^{*}}\right\}$. The TrimCaching Gen algorithm will repeat the above steps until all edge servers cannot cache more models. The proposed algorithm is outlined in Algorithm 4. The proposed algorithm can solve $\mathcal{P}1.1$ with the time complexity of $O\left(MI\right)$. In addition, we have the following result for the TrimCaching Gen algorithm.
@@ VII-A Simulation Setup
  • The special case: In the ResNet-based model library, the number of shared parameter blocks is independent of the scale of the model library because all of them are fine-tuned from three pre-trained models. Without loss of generality, we adopt the classic bottom layer freezing techniques to fine-tune these three models for the 100 downstream classification services. The number of frozen bottom layers for the ResNet18, ResNet34, and ResNet50 falls within the ranges of $\left[29,40\right]$, $\left[49,72\right]$, and $\left[87,106\right]$, respectively.
@@
TABLE II: Association Relationship of Superclasses
First-round FT Second-round FT
@@ VII-E Robustness to User Mobility
Fig. 11 shows the robustness of the TrimCaching Spec algorithm in the special case and that of the TrimCaching Gen algorithm in the general case. While the cache hit ratio degrades over time due to user mobility, the performance of the TrimCaching Spec algorithm and the TrimCaching Gen algorithm only degrades about by 3.91% and 2.83% over 2 h. This shows the effectiveness of the TrimCaching framework in the long run even though user mobility has not been explicitly considered in our problem formulation. This also implies that model replacement does not need to be re-conducted frequently, thereby saving backbone bandwidth resources.
@@ VIII Conclusions
In this paper, we proposed the TrimCaching framework for edge AI model caching. Specifically, by taking advantage of the shared parameter blocks among AI models for efficient storage, we investigated the parameter-sharing AI model placement problem. Considering multi-edge scenarios, the model placement problem was formulated to maximize the cache hit ratio under the constraints of model downloading latency and storage capacity. Then, we identified this problem as a submodular maximization problem with submodular constraints, revealing that no polynomial-time algorithm can solve it with a constant approximation ratio. To tackle this challenge, we first investigated the special case with a small fixed number of shared parameter blocks independent of the problem scale. We developed a polynomial-time algorithm with $\left(1-\epsilon\right)/2$-approximation guarantee. After that, we addressed the general case by devising a greedy algorithm that is also efficient and effective. It has been shown that the proposed TrimCaching framework significantly improves the cache hit ratio compared with traditional content placement strategies without considering shared parameter blocks among AI models. We hope this work will inspire further research on edge AI model caching, which has the potential to become a vital component in 6G edge intelligence.

### v2

@@ V-A Successive Greedy Algorithm and the Constant-approximation Guarantee
Suppose that each sub-problem can be solved optimally. In the considered special case, the TrimCaching Spec algorithm can obtain a solution $\hat{{\bf{X}}}$ for $\mathcal{P}1.1$ which is lower bounded by $U\left(\hat{{\bf{X}}}\right)\geq\frac{1}{2}U\left({\bf{X}}^{*}\right)$, where ${\bf{X}}^{*}$ is the optimal solution to $\mathcal{P}1.1$.
@@
Thus, a condition to achieve the constant-approximation guarantee of the TrimCaching Spec algorithm is obtaining the optimal solution to $\mathcal{P}2.1_{m}$. In the next subsection, we will introduce how to solve $\mathcal{P}2.1_{m}$ optimally in polynomial time.
@@ V-B DP-based Approach and the $\epsilon$-optimal Solution
We propose a DP-based algorithm to solve $\mathcal{P}2.1_{m}$. The submodularity of the constraint in $\mathcal{P}2.1_{m}$ results from the shared parameter blocks among models. To address the submodular constraints, our core idea is to make the caching decisions for shared parameter blocks and specific parameter blocks separately. We use $\mathcal{A}$ to denote a set where each element represents a combination of shared parameter blocks. All the possible combinations are included in $\mathcal{A}$. For example, $\{1,2,3,4,5,15\}$ and $\{10,12,13,15\}$ are two elements in $\mathcal{A}$ for Fig. 4. For a combination $\mathcal{N}$, all models that include $\mathcal{N}$ is denoted by $\mathcal{I}_{\mathcal{N}}$. For example, if $\mathcal{N}$ is $\{1,2,3,4,5,15\}$, $\mathcal{I}_{\mathcal{N}}$ is the set of model 1 and 2 for Fig. 4. Additionally, we use $d_{\mathcal{N}}$ to denote the occupied capacity of edge server $m$ for the shared parameter block combination $\mathcal{N}$. In the above example, $d_{\mathcal{N}}$ is the total size of parameter blocks 1, 2, 3, 4, 5, 15.
@@
The process of updating $\mathcal{T}\left(e_{\mathcal{N}},w_{\mathcal{N}}\right)$ for edge server $m$ is shown in Fig. 5. Specifically, on the one hand, if $w_{\mathcal{N}}\geq\frac{u\left(m,e_{\mathcal{N}}\right)}{\delta_{m,\mathcal{N}}}$, i.e., the number of cache hits of model $e_{\mathcal{N}}$ is smaller than the $w_{\mathcal{N}}$-th value of the number of cache hits, $\mathcal{T}\left(e_{\mathcal{N}},w_{\mathcal{N}}\right)$ is the minimum value between $\mathcal{T}\left(e_{\mathcal{N}}-1,w_{\mathcal{N}}\right)$ and $\mathcal{T}\left(e_{\mathcal{N}}-1,w_{\mathcal{N}}-\frac{u\left(m,e_{\mathcal{N}}\right)}{\delta_{m,\mathcal{N}}}\right)+D_{\mathcal{N}}\left(e_{\mathcal{N}}\right)$. If $\mathcal{T}\left(e_{\mathcal{N}}-1,w_{\mathcal{N}}\right)$ is smaller, it means that a combination of certain models in the first $e_{\mathcal{N}}-1$ models can achieve the $w_{\mathcal{N}}$-th value of the number of cache hits with a smaller data size. Thus, it is also the data size for the first $e_{\mathcal{N}}$ models to achieve the $w_{\mathcal{N}}$-th value of the number of cache hits. If $\mathcal{T}\left(e_{\mathcal{N}}-1,w_{\mathcal{N}}-\frac{u\left(m,e_{\mathcal{N}}\right)}{\delta_{m,\mathcal{N}}}\right)+D_{\mathcal{N}}\left(e_{\mathcal{N}}\right)$ is smaller, it represents the $w_{\mathcal{N}}$-th value of the number of cache hits $w_{\mathcal{N}}\delta_{m,\mathcal{N}}$ is the sum of the number of cache hits $w_{\mathcal{N}}\delta_{m,\mathcal{N}}-u\left(m,e_{\mathcal{N}}\right)$ of some models in the first $e_{\mathcal{N}}-1$ models and the number of cache hits $u\left(m,e_{\mathcal{N}}\right)$ of model $e_{\mathcal{N}}$, thus adding $D_{\mathcal{N}}\left(e_{\mathcal{N}}\right)$ into $\mathcal{T}\left(e_{\mathcal{N}}-1,w_{\mathcal{N}}-\frac{u\left(m,e_{\mathcal{N}}\right)}{\delta_{m,\mathcal{N}}}\right)$. On the other hand, if $w_{\mathcal{N}}<\frac{u\left(m,e_{\mathcal{N}}\right)}{\delta_{m,\mathcal{N}}}$, i.e., the number of cache hits of model $e_{\mathcal{N}}$ is already greater than the $w_{\mathcal{N}}$-th value of the number of cache hits, the $w_{\mathcal{N}}$-th value of the number of cache hits can only be realized by certain models in the first $e_{\mathcal{N}}-1$ models. Finally, after updating all $\mathcal{T}\left(e_{\mathcal{N}},w_{\mathcal{N}}\right)$, the maximum expected number of cache hits of edge server $m$ with $\mathcal{N}$, i.e., $w^{*}_{\mathcal{N}}\delta_{m,\mathcal{N}}$, can be determined, where
@@
where ^$\hat{\mathcal{I}}_{m}=\left\{i\mid\hat{x}_{m,i}\in\hat{{\bf{X}}}_{m}\right\}$. Thus, the optimal solution to $\mathcal{P}2.1_{m}$ is obtained.
@@
To further accelerate the computation of (18), we propose a rounding method to balance precision and computational efficiency. When determining the maximum expected number of cache hits of $\mathcal{I}_{\mathcal{N}}$, a smaller value of $\delta_{m,\mathcal{N}}$ leads to a larger value of $\frac{\sum\limits_{i\in\mathcal{I}_{\mathcal{N}}}u\left(m,i\right)}{\delta_{m,\mathcal{N}}}+1$ and a larger number of feasible $w_{\mathcal{N}}$, further increasing the complexity. To facilitate the execution, we define a factor $\epsilon$. Then, we round $u\left(m,i\right)$ into
@@
where $u_{m,\min}=\mathop{\min}\limits_{i\in\mathcal{I}}u\left(m,i\right)$. $\epsilon$ is a constant and $\epsilon\in\left(0,1\right]$. As a result, when determining the maximum expected cache hit ratio for edge server $m$, some $u\left(m,i\right)$ of models in $\mathcal{I}_{\mathcal{N}}$ can be rounded into the same value, and the number of feasible values of $w_{\mathcal{N}}$ can be decreased, thereby accelerating the DP process for $\mathcal{I}_{\mathcal{N}}$. In the rounding process, we denote $\dot{\delta}_{m,\mathcal{N}}$ by the granularity of $\dot{u}\left(m,i\right)$ for models in $\mathcal{I}_{\mathcal{N}}$. Similarly, $\mathcal{T}\left(e_{\mathcal{N}},\dot{w}_{\mathcal{N}}\right)$ is the smallest data size of specific parameter blocks that have to be cached on edge server $m$ for achieving the $\dot{w}_{\mathcal{N}}$-th value of the cache hits with the first $e_{\mathcal{N}}$ models in $\mathcal{I}_{\mathcal{N}}$. Besides, we define $\dot{w}^{*}_{\mathcal{N}}=\mathop{\arg\max}\limits_{\dot{w}_{\mathcal{N}}}\{\dot{w}_{\mathcal{N}}\mid\mathcal{T}\left(\left|\mathcal{I}_{\mathcal{N}}\right|,\dot{w}_{\mathcal{N}}\right)\leq Q_{m}-d_{\mathcal{N}}\}$ and $\dot{w}^{*}_{\mathcal{N}^{*}}\delta_{m,\mathcal{N}^{*}}=\mathop{\max}\limits_{\mathcal{N}}\{\dot{w}^{*}_{\mathcal{N}}\delta_{m,\mathcal{N}}\}$. At last, in $\mathcal{P}2.1_{m}$, the model caching decision corresponding to $\dot{w}^{*}_{\mathcal{N}^{*}}$ is denoted by $\dot{{\bf{X}}}_{m}$, and
@@
The TrimCaching Spec algorithm has a polynomial-time computational complexity $O\left(MI\right)$.
@@
The TrimCaching Spec algorithm can obtain a solution $\hat{{\bf{X}}}$ for $\mathcal{P}1.1$ satisfying $U\left(\hat{{\bf{X}}}\right)\geq\frac{1-\epsilon}{2}U\left({\bf{X}}^{*}\right)$.
@@ VI The General Case: Arbitrary Parameter Sharing
This section examines the general case of $\mathcal{P}1.1$, where models can arbitrarily share parameter blocks, in contrast to the special case with a small fixed number of shared parameter blocks. Formally speaking, the number of shared parameter blocks among models may increase with the scale of the model library, such that searching for all combinations of shared parameter blocks, as required by the TrimCaching Spec algorithm, results in an exponential time complexity in terms of the problem scale, which should be avoided in the general case.
Our proposed greedy algorithm for $\mathcal{P}1.1$ is outlined in Algorithm 4. The key part of the TrimCaching Gen algorithm is in line 4. In the $l$-th step, the algorithm first calculates the incremental cache hit ratio for each $x_{m,i}$ based on the previous placement decision ${\bf{X}}^{l-1}$ made in the $l-1$-th step. Subsequently, the TrimCaching Gen algorithm identifies $\{m^{*},i^{*}\}$ which yields the maximum increased cache hit ratio, i.e., $U\left({\bf{X}}^{l-1}\cup\left\{x_{m^{*},i^{*}}\right\}\right)-U\left({\bf{X}}^{l-1}\right)$, while ensuring that $g_{m}\left({\bf{X}}_{m}^{l-1}\cup\{x_{m^{*},i^{*}}\}\right)$ does not exceed the edge server capacity $Q_{m}$, where ${\bf{X}}_{m}^{l-1}$ is the model placement result of edge server $m$ in the $l-1$-th step. By adding $x_{m^{*},i^{*}}$ into ${\bf{X}}^{l-1}$, we have ${\bf{X}}^{l}={\bf{X}}^{l-1}\cup\left\{x_{m^{*},i^{*}}\right\}$. The TrimCaching Gen algorithm will repeat the above steps until all edge servers cannot cache any models. The proposed algorithm can solve $\mathcal{P}1.1$ with the time complexity of $O\left(MI\right)$. In addition, we have the following result for the TrimCaching Gen algorithm.
@@ VII-A Simulation Setup
  • The special case: In the ResNet-based model library, the number of shared parameter blocks is independent of the scale of the model library because all of them are fine-tuned from three pre-trained models. Without loss of generality, we adopt the classic bottom layer freezing techniques to fine-tune these three models for the 100 downstream classification services. The number of frozen bottom layers (each corresponding to a shared parameter block in our system model) for ResNet18, ResNet34, and ResNet50 falls within the ranges of $\left[29,40\right]$, $\left[49,72\right]$, and $\left[87,106\right]$, respectively.
@@
TABLE II: Fine-tuning Settings
First-round fine-tuning Second-round fine-tuning
@@ VII-E Robustness to User Mobility
Fig. 11 shows the robustness of the TrimCaching Spec algorithm and the TrimCaching Gen algorithm for the special case in the GPT2-based model library. While the cache hit ratio degrades over time due to user mobility, the performance of the TrimCaching Spec algorithm and the TrimCaching Gen algorithm only degrades by about 3.91% and 2.83% over 2 h. This shows the effectiveness of the TrimCaching framework in the long run, even though user mobility has not been explicitly considered in our problem formulation. This also implies that model replacement does not need to be re-conducted frequently, thereby saving backbone bandwidth resources.
@@ VIII Conclusions
In this paper, we proposed the TrimCaching framework for edge AI model caching. Specifically, by taking advantage of the shared parameter blocks among AI models for efficient storage, we investigated the parameter-sharing AI model placement problem. Considering multi-edge scenarios, the model placement problem was formulated to maximize the cache hit ratio under the constraints of service latency requirement and storage capacity. Then, we identified this problem as a submodular maximization problem with submodular constraints, revealing that no polynomial-time algorithm can solve it with a constant approximation ratio. To tackle this challenge, we first investigated the special case with a small fixed number of shared parameter blocks independent of the problem scale. We developed a polynomial-time algorithm with $\left(1-\epsilon\right)/2$-approximation guarantee. After that, we addressed the general case by devising a greedy algorithm that is also efficient and effective. It has been shown that the proposed TrimCaching framework significantly improves the cache hit ratio compared with traditional content placement strategies without considering shared parameter blocks among AI models. We hope this work will inspire further research on edge AI model caching, which has the potential to become a vital component in 6G edge intelligence.